# Development of a Neutron Imager using CMOS Pixelated Sensor

### v1

@@ Abstract
We acquired an image of the Siemens Star Chart using the 10B-INTPIX4 CMOS pixelated imager at J-PARC MLF BL21 (NOVA). The image blurriness, characterized by a root mean square (rms) value of 17 $\mu$m in the line spread function, corresponds to an rms beam divergence of 2.7 mrad. for the neutron beam. This is equivalent to the L/D ratio of 370. The contrast of the image was analyzed as a function of structural scale of the chart, and based on this contrast, the modulation transfer function (MTF) of the imaging system was evaluated. The results indicate that the structure corresponding to the 50% cutoff is 12 line-space pairs per mm.
neutron imaging, monolithic CMOS sensor, spacial resolution, modulation transfer function
@@ 1 Introduction
In this paper, we use the SOIPIX-INTPIX4 monolithic CMOS pixel sensor [2] and present the results of imaging a Siemens Star Chart, a standard tool for evaluating imaging performance, at the J-PARC, Materials and Life Science Experimental Facility (MLF) – Beam Line 21 (BL21/NOVA) [3]. By analyzing the contrast dependence on the line-space pitch, we evaluated the modulation transfer function (MTF) of the imaging system.
The structure of this paper is as follows: Chapter 2 introduces the experimental components, Chapter 3 discusses the selection of neutron events, and Chapter 4 examines the imaging accuracy. Finally, Chapter 5 presents the summary.
@@ 2 Experimental Components
Previously, an imaging experiment was conducted at J-PARC MLF BL10 (NOBORU) [7], where a Gd-contained line-space pattern was captured and analyzed. The results confirmed that the spatial resolution, as evaluated by the root mean square (rms) of the Line Spread Function (LSF), was below 5 $\mu$m [1]. More recently, simultaneous detection of charged particle pairs from the converter had been demonstrated by sandwiching it between two sensors [8]. The sensor’s response to neutrons can be regarded well understood.
@@
10B-INTPIX4 was placed 6.5 m downstream from the NOVA’s standard sample location, at a distance of 21 m from the neutron production target with a moderator, just before the neutron beam dump duct. The neutrons primarily travel through a vacuum duct, with the exception of a section totaling approximately 3.5 m (3.1 m upstream and 0.4 m downstream of the sample chamber) where they pass through air. The neutron beam provided covers a wavelength range from 0.12 Å to 8.3 Å and can be considered to behave as a quantum beam for our device. On the sensor surface, the beam had an approximate size of $70\times 70$ mm2, illuminating the entire sensor area. Active control of beam divergence was not implemented.
@@ 3 Event Identification
The figure below in the Fig. 2 is a scatter plot of the $y$-component of $\vec{m}_{2}$ versus $m_{0}$. Neutron-induced events are distributed along the dashed line (called it as a neutron line, hereafter). Events around $m_{0}=2200$ channels originate from 7Li particles, while those around $m_{0}=5500$ channels originate from $\alpha$ particles. Events falling within $|\Delta m_{2,y}|<0.18$ pixel2 along the neutron line are extracted as neutron events, where $\Delta m_{2,y}$ represents the deviation from it. The upper graph in the Fig. 2 shows the $m_{0}$ distribution of these extracted neutron events. The low-energy side includes contributions from $\gamma$-ray background, while the high-energy side consists of multiple-count events where multiple event clusters are present within the $7\times 7$ cluster region. To eliminate these contributions, the range of 1400 channels $<m_{0}<$ 7000 channels is utilized for the subsequent analysis.
Figure 3 illustrates the Siemens Star Chart image obtained from the experiment. This figure is presented as a two-dimensional histogram with square bins of 0.2 pixels in width, with the number of entries per bin indicated by the accompanying color bar. Approximately 10 events par bin were recorded. The areas with fewer events near the edges of the sensor correlate with regions where the 10B film is thinner.
Figure 3: Obtained image of the Siemens Star Chart. The spokes in the outer ring of the first doughnut shape are clearly distinguishable. Toward the innermost ring, the chart image gradually becomes unclear. This is consistent with the effects of beam divergence.
@@ 4 Imaging Accuracy
The sharpness of the obtained images is influenced by the inherent spatial resolution of the imager, the accuracy of the evaluation chart, and the divergence of the beam. In our setup, the chart’s accuracy is sufficiently high to be considered negligible, and with the imager’s spatial resolution below 5 $\mu$m, its impact is expected to be minimal. If the image blurriness is primarily attributed to beam divergence, given that the distance between the chart and the sensor is 6 mm, the estimated beam divergence is 2.7 mrad. This corresponds to a neutron source size of 60 mm and the L/D ratio of 370, which aligns well with the NOVA design’s L/D ratio of approximately 230.
@@ 5 Summary
At J-PARC MLF BL21 (NOVA), image of the Siemens Star Chart was obtained using the 10B-INTPIX4 neutron imager. Upon analyzing the image, it is determined that the sharpness of the obtained image corresponds to a blurriness of 1.0 pixel $=17$ $\mu$m, as indicated by the rms of the LSF. If this blurriness is attributed to beam divergence, it corresponds to an rms beam divergence of 2.7 mrad. suggesting that the effective size of the neutron source is approximately 60 mm and the L/D ratio of 370. This device may potentially be utilized to estimate the size of neutron sources that cannot be directly observed. The structure corresponding to the 50% cutoff of the MTF is found to be 12 line-space pairs per mm.

### v2

@@ Abstract
We acquired an image of the Siemens Star Chart using the 10B-INTPIX4 CMOS pixelated imager at J-PARC MLF BL21 (NOVA). The image blurriness, characterized by a root mean square (rms) value of 17 $\mu$m in the line spread function, corresponds to an rms beam divergence of 2.7 mrad. for the neutron beam. This is equivalent to the effective geometrical ratio (L/D ratio) of 370. The contrast of the image was analyzed as a function of structural scale of the chart, and based on this contrast, the modulation transfer function (MTF) of the imaging system was evaluated. The results indicate that the structure corresponding to the 50% cutoff is 12 line-space pairs per mm.
@@ 1 Introduction
In this paper, we use the SOIPIX-INTPIX4 monolithic CMOS pixel sensor [2] and present the results of imaging a Siemens Star Chart, a standard tool for evaluating imaging performance, at the J-PARC, Materials and Life Science Experimental Facility (MLF) – Beam Line 21 (BL21/NOVA) [3]. To understand the capabilities of the high-precision imaging system, we evaluated the image blurring using multiple metrics, including the root mean square (rms) in the line spread function (LSF), the modulation transfer function (MTF), the 50% cutoff line-space pair width, and the effective geometrical ratio (L/D ratio).
The structure of this paper is as follows: Section 2 introduces the experimental components, Section 3 discusses the selection of neutron events, and Section 4 examines the imaging accuracy. Finally, Section 5 presents the summary.
@@ 2 Experimental Components
Previously, an imaging experiment was conducted at J-PARC MLF BL10 (NOBORU) [7], where a Gd-contained line-space pattern was captured and analyzed. The results confirmed that the spatial resolution, as evaluated by the rms of the LSF, was below 5 $\mu$m [1]. More recently, simultaneous detection of charged particle pairs from the converter had been demonstrated by sandwiching it between two sensors [8]. The sensor’s response to neutrons can be regarded well understood.
@@
10B-INTPIX4 was placed 6.1 m downstream from the NOVA’s standard sample location, at a distance of 21 m from the neutron production target with a moderator, just before the neutron beam dump duct. The neutrons primarily travel through a vacuum duct, with the exception of sections totalling approximately 3.5 m (3.1 m upstream and 0.4 m downstream of the sample chamber) where they pass through air. The neutron beam provided covers a wavelength range from 0.12 Å to 8.3 Å and can be considered to behave as a quantum beam for our device. On the sensor surface, the beam had an approximate size of more than $40\times 40$ mm2, illuminating the entire sensor area. A 20 mm $\times$ 20 mm collimator is located 6.5 meters upstream of the imager, determining the geometrical L/D ratio of the beamline to be 330.
@@ 3 Event Identification
The figure below in the Fig. 2 is a scatter plot of the $y$-component of $\vec{m}_{2}$ versus $m_{0}$. Neutron-induced events are distributed along the dashed line (called it as a neutron line, hereafter). Events around $m_{0}=2200$ channels originate from 7Li particles, while those around $m_{0}=5500$ channels originate from $\alpha$ particles. Events falling within $|\Delta m_{2,y}|<0.18$ pixel2 along the neutron line are extracted as neutron events, where $\Delta m_{2,y}$ represents the deviation from it. The upper graph in the Fig. 2 shows the $m_{0}$ distribution of these extracted neutron events. The low-energy side includes contributions from $\gamma$-ray background, while the high-energy side consists of multiple-count events where multiple event clusters are present within the $7\times 7$ pixel cluster region. To eliminate these contributions, the range of 1400 channels $<m_{0}<$ 7000 channels is utilized for the subsequent analysis.
Figure 3 illustrates the Siemens Star Chart image obtained from the experiment. This figure is presented as a two-dimensional histogram with square bins of 0.2 pixels in width, with the number of entries per bin indicated by the accompanying color bar. Approximately 10 events per bin were recorded. The areas with fewer events near the edges of the sensor correlate with regions where the 10B film is thinner.
Figure 3: Obtained image of the Siemens Star Chart. The spokes in the outer ring of the first doughnut shape are clearly distinguishable. Toward the innermost ring, the chart image gradually becomes unclear due to the effect of beam divergence.
@@ 4 Imaging Accuracy
The sharpness of the obtained images is influenced by the inherent spatial resolution of the imager, the accuracy of the evaluation chart, and the divergence of the beam. In our setup, the chart’s accuracy is sufficiently high to be considered negligible, and with the imager’s spatial resolution below 5 $\mu$m, its impact is expected to be minimal. If the image blurriness is primarily attributed to beam divergence, given that the distance between the chart and the sensor is 6 mm, the estimated beam divergence is 2.7 mrad. This corresponds to an effective size of the neutron beam at the collimator is approximately 18 mm, and the L/D ratio of 370, which aligns well with the NOVA design’s geometrical L/D ratio of approximately 330.
@@ 5 Summary
At J-PARC MLF BL21 (NOVA), image of the Siemens Star Chart was obtained using the 10B-INTPIX4 neutron imager. Upon analyzing the image, it is determined that the sharpness of the obtained image corresponds to a blurriness of 1.0 pixel $=17$ $\mu$m, as indicated by the rms of the LSF. If this blurriness is attributed to beam divergence, it corresponds to an rms beam divergence of 2.7 mrad. suggesting that the effective size of the neutron beam at the collimator is approximately 18 mm and the L/D ratio of 370. By configuring the system to capture the entire target plane, this device can be effectively used to estimate or monitor the size of a neutron beam, even at locations such as neutron production targets that are inaccessible due to high radiation levels. The structure corresponding to the 50% cutoff of the MTF is found to be 12 line-space pairs per mm.